A luminaire is mounted 2.8m above the floor and it is required to calculate the illuminance due to the luminaire at a point on the floor that is 1.5m from the point directly beneath the luminaire.
What is the hypotenuse of the triangle formed by the luminaire and the two points on the floor?
and what are the two angles formed by the above triangle that are not right angles?

Answer by Cromlix(4381)

Hi there,
Hypotenuse^2 = 2.8^2 + 1.5^2
Hypotenuse^2 = 10.09
Hypotenuse = √10.09
Hypotenuse = 3.2 m
Angles:
tan(x) = 1.5/2.8
x = 28.2 degrees
tan(x) = 2.8/1.5
x = 61.8 degrees
Hope this helps :-)